# Semantic Reasoning Meets Numerical Precision: An LLM-Powered Multi-Agent System for Power Grid Control

### v1

@@ Abstract
The increasing penetration of Distributed Energy Resources (DERs), widespread adoption of Electric Vehicles (EVs), and the growing frequency of extreme weather events have significantly increased the complexity of power grid planning, operation, and management. Traditional rule-based systems and numerical optimization approaches often struggle with the scale, dynamics, and adaptability required by modern power networks. This paper introduces Grid-Agent, an autonomous, AI-driven framework that combines Large Language Models (LLMs) with multi-agent reinforcement learning to detect and remediate grid violations in real time. Grid-Agent integrates semantic reasoning with numerical precision through a modular agent architecture: a planning agent generates coordinated action sequences using numerical power flow solvers, while a validation agent evaluates system stability and action effectiveness via sandboxed execution with safety rollbacks. To ensure scalability, Grid-Agent incorporates an adaptive multiscale network representation that dynamically selects optimal encoding schemes based on network size and complexity. The framework enables coordinated violation resolution through optimizing switch configurations, battery deployment, and load curtailment strategies. Experimental results in standard IEEE and CIGRE test systems (IEEE 69-bus, CIGRE MV, and IEEE 30-bus) demonstrate superior violation mitigation performance. Additionally, the framework’s built-in data collection and learning capabilities enable continuous learning and adaptation to diverse network topologies. The autonomous nature of the framework makes it particularly suitable for modern smart grid applications requiring rapid response to dynamic operating conditions.
@@ I Introduction
Modern power grids face unprecedented challenges due to increased integration of renewable energy, electric vehicles (EVs), increasing frequency of extreme weather events, and aging infrastructure. These factors demand optimization strategies that are not only intelligent and responsive, but also adaptable to highly dynamic and uncertain operating conditions, especially as grid failures and blackouts have been estimated to cost the U.S. economy over \$150 billion annually . Traditional power system optimization methods are based on fixed rule-based logic and deterministic numerical algorithms, often fall short in their ability to generalize, interpret contextual information, or reason semantically—capabilities that are essential for managing the complexity of real-world grid scenarios.
The emergence of Large Language Models (LLMs) presents transformative opportunities for power system applications by enabling the integration of natural language reasoning with numerical precision. Recent advancements in LLM-based multi-agent systems have demonstrated impressive capabilities in solving complex, multi-faceted problems across various domains [1]. However, their application to mission-critical infrastructures such as electric power grids remains largely unexplored.
This paper proposes Grid-Agent, an innovative multi-agent framework that harnesses the power of LLMs to autonomously interpret network topology, detect and resolve electrical violations, and help system recovery from events such as cyber-attacks or blackouts. Unlike conventional approaches that treat power system optimization as isolated numerical problems, Grid-Agent integrates semantic understanding with domain-specific analysis tools to deliver coordinated, interpretable, and explainable solutions.
@@
  • Adaptive Multiscale Representation. A dynamic encoding scheme was developed that automatically adjusts network representation granularity based on system size and complexity, ensuring scalability from microgrids to large distribution networks.
@@
  • Continuous Learning Capability. A training data generation framework that creates high-quality datasets in multiple formats for LLM fine-tuning to enable the system to learn from operational experience and improve performance over time.
The remainder of this paper is organized as follows: Section II reviews related work in power systems optimization and LLM-based multi-agent systems. Section III presents the Grid-Agent system architecture and methodology. Section IV describes the experimental setup and validation. Section V discusses results and performance analysis. Section VI concludes with future research directions.
@@ II-A Power Systems Optimization
Traditional power system optimization has progressed from manual control to sophisticated automated frameworks. Classical methods, such as Optimal Power Flow (OPF), have been widely applied to objectives like peak shaving, cost/loss minimization, and violation mitigation. While mathematically robust, these approaches often face convergence issues due to the complexity of modern grids, which incorporate high levels of distributed renewable energy, highly variable loads, and numerous discrete controllable assets such as switches and tap changers. These factors result in a mixed-integer, non-convex, nonlinear, and unbalanced AC optimal powerflow problem, which is NP-hard and computationally intensive to solve. Furthermore, their rigid mathematical structure makes it difficult to incorporate heuristic operator knowledge or adapt to unforeseen event types.
@@ II-B Large Language Models in Engineering Applications
In the energy sector, preliminary investigations have examined LLMs for grid analytics [5], energy trading [6], and demand response optimization [7]. However, these applications primarily focus on data analysis and prediction rather than direct system control and optimization. This leaves a critical gap in leveraging LLM reasoning for active, real-time control and optimization in critical infrastructure, which is the primary focus of our work.
@@ III-A Power Grid Violation Resolution Problem
A violation $\mathcal{V}$ occurs when the system state breaches predefined operational limits. We consider three primary categories of violations:
  • Voltage Violations ($\mathcal{V}_{\text{volt}}$): Occur when a bus voltage magnitude $V_{i}$ falls outside its acceptable range $V_{\min},V_{\max}]$.
    $\mathcal{V}_{\text{volt}}=\{i\in{N}\mid V_{i}<V_{\min}\text{ or }V_{i}>V_{\max}\}$ (1)
  • Thermal Violations ($\mathcal{V}_{\text{therm}}$): Arise when the apparent power flow $S_{ij}$ or current $I_{ij}$ on a line exceeds its maximum rating.
    $\mathcal{V}_{\text{therm}}=\{(i,j)\in{E}\mid S_{ij}>S_{ij}^{\max}\text{ or }I_{ij}>I_{ij}^{\max}\}$ (2)
  • Disconnected Buses ($\mathcal{V}_{\text{disc}}$): Happen when a bus becomes electrically isolated from the main grid, often resulting in a complete loss of power.
The core problem is to find an optimal sequence of control actions, ${A}=\{a_{1},a_{2},\dots,a_{k}\}$, that resolves all existing violations while minimizing operational costs and the complexity of the intervention. This can be formally expressed as an optimization problem:
$\min_{{A}}\sum_{t=1}^{k}c(a_{t})+\lambda|{A}|$ (3)
subject to the constraints defined in the following subsections. Here, $c(a_{t})$ is the cost associated with control action $a_{t}$, |A||{A}| is the total number of actions, and $\lambda$ is a penalty parameter that encourages solution parsimony. In essence, the objective is to find the most economical and least disruptive solution to restore the grid to a secure state.
@@ III-C Action Space Constraints
The control actions available to Grid-Agent are constrained by the physical and operational capabilities of the grid equipment. The action space includes:
  • Switch Operations: Binary decisions to open or close a switchable line $(i,j)$ within the set of designated switchable lines ${E}_{\text{sw}}\subseteq{E}$.
    $s_{ij}\in\{0,1\},\quad\forall(i,j)\in{E}_{\text{sw}}$ (6)
  • Load Curtailment: Continuous reduction of the power demand $P_{i}^{\text{load}}$ at buses within the set of curtailable loads ${N}_{\text{L}}\subseteq{N}$. The reduction is determined by a curtailment fraction $\gamma_{i}$.
    $P_{i}^{\text{new\_load}}=P_{i}^{\text{orig\_load}}\cdot(1-\gamma_{i})$ (7)
    $0\leq\gamma_{i}\leq\gamma_{\max},\quad\forall i\in{N}_{\text{L}}$ (8)
  • Battery Placement and Dispatch: This involves two coupled decisions: (1) a discrete decision on where to place a battery, and (2) a continuous decision on how much active and reactive power it should inject or absorb.
    The placement is a binary decision $b_{i}$ for each bus, constrained by the total number of available batteries, $B_{\text{total}}$.
    $\sum_{i\in{N}}b_{i}\leq B_{\text{total}},\quad\text{with }b_{i}\in\{0,1\}$ (9)
    The dispatch consists of setting the active power $P_{B,i}$ and reactive power $Q_{B,i}$ output of the battery at bus $i$. These outputs are constrained by the battery’s maximum power rating, $S_{B}^{\max}$. Critically, power can only be dispatched if a battery is placed at that bus ($b_{i}=1$). This coupling is enforced as follows:
    $P_{B,i}^{2}+Q_{B,i}^{2}\leq(b_{i}\cdot S_{B}^{\max})^{2}$ (10)
    $-P_{B}^{\max}\leq P_{B,i}\leq P_{B}^{\max}$ (11)
    $-Q_{B}^{\max}\leq Q_{B,i}\leq Q_{B}^{\max}$ (12)
    This formulation couples the discrete placement decision with the continuous operational control, ensuring a battery is both placed and operated within its physical limits to support the grid
@@ III-D1 LLMs as Reasoning Engines for Grid Management
Modern LLMs function as powerful reasoning engines. In the context of power systems, they offer several key advantages over traditional algorithms:
  • Semantic Understanding: LLMs can process and interpret not just raw numerical data, but also heterogeneous information such as network topology descriptions, operational logs, and maintenance reports. This allows for a more holistic understanding of the state of the grid.
  • Contextual Reasoning: They can incorporate high-level domain knowledge and operational constraints (e.g., ”prioritize critical loads”, ” avoid switching near a known weak point”) that are difficult to encode in conventional optimization models.
  • Tool-Augmented Operation: Crucially, recent advancements allow LLMs to use external tools. For power systems, this means an LLM can delegate complex numerical calculations to a validated power flow solver, grounding its high-level reasoning in the physical realities and precise constraints of the grid.
@@ III-D2 Multi-Agent Systems for Decomposed Problem Solving
A single monolithic system struggles to manage the entire scope of violation resolution. Agentic AI provides a modular and scalable solution by decomposing the problem into specialized sub-tasks, each handled by a dedicated agent. This paradigm offers significant benefits:
  • Specialized Expertise: Different agents can be developed to excel at specific functions, such as topology analysis, action planning, solution validation, or results summarization.
  • Iterative Refinement: Agents can collaborate in a loop, progressively improving solutions. For example, a Planner agent can propose an action, and a Validator agent can simulate its impact, providing feedback that allows the Planner to refine its strategy.
  • Scalability and Interpretability: The modular nature allows the system to adapt to grids of varying complexity, while the explicit interactions between agents create a transparent and auditable reasoning chain for every decision.
The Multi-Agent System can be formulated as:
$W=(A,F,C)$
where $W$ represents the workflow, $A=a_{1},a_{2},...,a_{m}$ is the set of agents, $F$ defines the flow relationships between agents, and $C$ specifies conditional branching logic.
The workflow state $S_{t}$ at iteration $t$ captures the complete system state, including network configuration, detected violations, executed actions, and intermediate results. State transitions are governed by:
$S_{t+1}=\mathcal{T}(S_{t},a_{i},\theta)$
where $\mathcal{T}$ is the transition function, $a_{i}$ is the executing agent, and $\theta$ represents the agent’s parameters and decision logic.
This formulation enables systematic exploration of the action space while maintaining solution quality and computational efficiency through intelligent agent coordination.
@@ IV System Architecture and Methodology
The Grid-Agent framework is an autonomous multi-agent system designed to resolve power grid violations by synergistically integrating the semantic reasoning of Large Language Models (LLMs) with the numerical precision of validated power system simulations. The methodology is founded on a state-driven workflow where specialized agents collaborate to plan, execute, and validate control actions in a secure, sand-boxed environment. This section details the four core pillars of this methodology: the multi-agent workflow, the LLM-guided action planning process, the adaptive network representation scheme, and the integrated safety mechanisms.
@@ IV-A State-Driven Multi-Agent Workflow
The framework’s operation is governed by a cyclic, state-driven workflow, depicted in Fig. 1. This workflow is orchestrated by five specialized agents, each with a distinct role:
  • Topology Agent: Initializes the resolution process. It parses the grid’s topology and operational data to establish the initial network state and identify the set of existing violations. This agent provides the foundational context for the planning cycle.
  • Planner Agent: Serves as the core reasoning engine. It receives the current network state and violation report from the central state object. Leveraging its LLM-based intelligence, it formulates a strategic, multi-step action plan designed to cohesively resolve the identified issues.
  • Executor Agent: Acts as the interface between strategic planning and numerical simulation. It translates the Planner’s abstract directives (e.g., ”reconfigure switches to isolate fault”) into concrete, tool-specific API calls for a numerical power flow solver. Critically, all actions are performed on a sand-boxed copy of the network to prevent any impact on the live system.
  • Validator Agent: Ensures the efficacy and safety of the proposed plan. After the Executor applies the actions, the Validator runs a new power flow analysis on the sand-boxed network. It assesses whether violations have been resolved without introducing new instabilities. Based on this outcome, it either approves the changes and terminates the workflow or signals the need for a new planning cycle. Its integrated rollback mechanism reverts any changes that fail to improve the network state, ensuring monotonic progress.
  • Summarizer Agent: Facilitates explainability and continuous learning. Upon successful resolution, this agent generates a human-readable explanation of the final solution and structures the entire process, initial state, actions, and final state, into a data entry for the continuous learning dataset.
@@ IV-B LLM-Guided Coordinated Action Planning
The central task of finding an optimal action sequence A (as formulated in Eq.(3)) is navigated by the Planner Agent. Unlike traditional optimizers, Grid-Agent uses the LLM as a tool-augmented reasoning engine. This is achieved through a comprehensive system prompt that structures the LLM’s task, defining five key elements:
@@
  1. 1.
    Role Definition: The LLM is instructed to act as an expert power system operator, grounding its responses in domain-specific best practices.
  2. 2.
    State Context: It receives a structured representation of the current network state, including topology, component statuses, and a detailed list of all active violations.
  3. 3.
    Action Space and Constraints: The prompt explicitly defines the set of available tools (e.g., update_switch_status, add_battery, curtail_load) and their associated operational constraints (e.g., maximum number of deployable batteries, available curtailable loads).
  4. 4.
    Strategic Guidance: The LLM is guided by a prioritized action policy that mirrors industry best practices: (1) Topology Reconfiguration, (2) Battery Deployment/Dispatch, and (3) Demand Response. This constrains the search space and aligns the LLM’s reasoning with established operational logic.
  5. 5.
    Output Schema: To ensure reliable machine integration, the LLM is required to respond exclusively with a machine-parsable list of tool calls in a structured format (e.g., JSON). This eliminates ambiguous natural language and enables direct interfacing with the Executor Agent.
IV-C Adaptive Multi-Scale Network Representation
IV-D Safety, Validation, and Rollback Mechanisms
Operational safety in a critical infrastructure context is paramount. Grid-Agent incorporates a multi-layered safety and validation strategy to ensure system stability and reliability:
  • Proactive Execution Check (Executor): Before applying any action in the sandbox, the Executor Agent performs a preliminary validation. It uses the power flow solver to ensure that the proposed API call is syntactically correct and will not cause an immediate simulation convergence failure. This acts as a first line of defense against ill-formed commands.
  • Post-Hoc State Assessment (Validator): After an action sequence is executed, the Validator Agent conducts a comprehensive assessment. It verifies not only that the target violations are resolved but also that no new violations (e.g., overloads, voltage deviations) have been introduced elsewhere in the network.
  • Monotonic Progress Guarantee (Rollback): If the Validator determines that the executed plan has failed to improve the grid state or has worsened it, the rollback mechanism is automatically triggered. It reverts the sand-boxed network to its state before the failed attempt, ensuring that the system only progresses towards a verifiably better solution.
This defense-in-depth approach, combining proactive checks with comprehensive post-hoc validation and automated rollbacks, provides the essential safety guarantees for deploying an AI-driven system in a mission-critical environment.
IV-E Continuous Learning Framework

### v2

# Grid-Agent: An LLM-Powered Multi-Agent System for Power Grid Control

@@ Abstract
Modern power grids face unprecedented complexity from the increasing penetration of Distributed Energy Resources (DERs), the widespread adoption of Electric Vehicles (EVs), and the growing frequency of extreme weather events. This complexity challenges traditional approaches to grid planning, operation, and management. Traditional rule-based systems and numerical optimization approaches often struggle with the scale, dynamics, and adaptability required by modern power networks. This paper introduces Grid-Agent, an autonomous, AI-driven framework that combines Large Language Models (LLMs) within a multi-agent framework to detect and remediate grid violations. Grid-Agent integrates semantic reasoning with numerical precision through a modular agent architecture: a planning agent generates coordinated action sequences using numerical power flow solvers, while a validation agent evaluates system stability and action effectiveness via sandboxed execution with safety rollbacks. To ensure scalability, Grid-Agent incorporates an adaptive multi-scale network representation that dynamically selects optimal encoding schemes based on network size and complexity. The framework enables coordinated violation resolution through optimizing switch configurations, battery deployment, and load curtailment strategies. Experimental results in standard IEEE and CIGRE test systems (IEEE 69-bus, CIGRE MV, and IEEE 30-bus) demonstrate superior violation mitigation performance. Additionally, the framework’s built-in data collection and learning capabilities enable continuous learning and adaptation to diverse network topologies. The autonomous nature of the framework makes it particularly suitable for modern smart grid applications requiring rapid response to dynamic operating conditions.
@@ I Introduction
Modern power grids face unprecedented challenges due to increased integration of renewable energy, electric vehicles (EVs), increasing frequency of extreme weather events, and aging infrastructure. These factors demand optimization strategies that are not only intelligent and responsive, but also adaptable to highly dynamic and uncertain operating conditions, especially as grid failures and blackouts have been estimated to cost the U.S. economy over \$150 billion annually [1]. Traditional power system optimization methods are based on fixed rule-based logic and deterministic numerical algorithms, often fall short in their ability to generalize, interpret contextual information, or reason semantically—capabilities that are essential for managing the complexity of real-world grid scenarios.
The emergence of Large Language Models (LLMs) presents transformative opportunities for power system applications by enabling the integration of natural language reasoning with numerical precision. Recent advancements in LLM-based multi-agent systems have demonstrated impressive capabilities in solving complex, multi-faceted problems across various domains [2] [3]. However, their application to mission-critical infrastructures such as electric power grids remains largely unexplored.
This paper introduces Grid-Agent, an autonomous, AI-driven framework conceived as an exploration into how LLMs can be applied to power system control, rather than as a finalized, production-ready solution. The primary goal is to push the boundaries of what LLMs and multi-agent systems might bring to future smart grid operations by demonstrating a novel methodology for autonomous grid management. Consequently, this work prioritizes the exploration of agentic reasoning, coordinated planning, and semantic interpretation over the formal optimality of the resulting control actions. The focus is on demonstrating a paradigm where Grid-Agent can interpret network topology, detect electrical violations, and formulate remediation strategies by integrating semantic understanding with domain-specific analysis tools.
@@
  • Adaptive Multi-scale Representation. A dynamic encoding scheme was developed that automatically adjusts network representation granularity based on system size and complexity, ensuring scalability from micro-grids to large distribution networks.
@@
  • Continuous Learning Capability. A training data generation framework that creates high-quality datasets in multiple formats for LLM fine-tuning to enable the system to learn from operational experience, ensuring its performance improves and adapts over time.
The remainder of this paper is organized as follows: Section II reviews related work in power systems optimization and LLM-based multi-agent systems. Section III formulates the power grid violation resolution problem. Section IV presents the Grid-Agent system architecture and methodology. Section V describes the experimental setup and validation. Section VI discusses results and future works. Section VII concludes the achievements.
@@ II-A Power Systems Optimization
Traditional power system optimization has progressed from manual control to sophisticated automated frameworks. Classical methods, such as Optimal Power Flow (OPF) [4], [5], have been widely applied to objectives like peak shaving, cost/loss minimization, and violation mitigation. While mathematically robust, these approaches often face convergence issues due to the complexity of modern grids, which incorporate high levels of distributed renewable energy, highly variable loads, and numerous discrete controllable assets such as switches and tap changers. These factors result in a mixed-integer, non-convex, nonlinear, and unbalanced AC optimal powerflow problem, which is NP-hard and computationally intensive to solve. Furthermore, their rigid mathematical structure makes it difficult to incorporate heuristic operator knowledge or adapt to unforeseen event types. Furthermore, these approaches often address isolated tasks and lack a unified framework for holistic, system-level optimization and control. This highlights the need for a system that can integrate high-level strategic reasoning with precise numerical validation, a core principle of our proposed Grid-Agent framework.
@@ II-B Large Language Models in Engineering Applications
In the energy sector, preliminary investigations have examined LLMs for grid analytics [12], energy trading [13], and demand response optimization [14]. However, these applications primarily focus on data analysis and prediction rather than direct system control and optimization. This leaves a critical gap in utilizing LLM reasoning for active, real-time control and optimization in critical infrastructure, which is the primary focus of our work.
@@ III-A Power Grid Violation Resolution Problem
A violation $\mathcal{V}$ occurs when the system state breaches predefined operational limits. We consider three primary categories of violations: voltage violations, thermal violations, and disconnected buses.
A voltage violation $\mathcal{V}_{\text{volt}}$ occurs when a bus voltage magnitude $V_{i}$ falls outside its acceptable range $V_{\min},V_{\max}]$. Formally, the set of buses experiencing voltage violations is defined as
$\mathcal{V}_{\text{volt}}=\{i\in N\mid V_{i}<V_{\min}\text{ or }V_{i}>V_{\max}\}.$ (1)
A thermal violation $\mathcal{V}_{\text{therm}}$ arises when the apparent power flow $S_{ij}$ or current $I_{ij}$ on a line exceeds its maximum rating. This can be expressed as
$\mathcal{V}_{\text{therm}}=\{(i,j)\in E\mid S_{ij}>S_{ij}^{\max}\text{ or }I_{ij}>I_{ij}^{\max}\}.$ (2)
Finally, a disconnected bus violation $\mathcal{V}_{\text{disc}}$ occurs when a bus becomes electrically isolated from the main grid, often resulting in a complete loss of power supply.
@@ III-C Action Space Constraints
The control actions available to Grid-Agent are constrained by the physical and operational capabilities of the grid equipment. The action space consists of switch operations, load curtailment, and battery placement with dispatch.
Switch operations are represented as binary decisions to open or close a switchable line $(i,j)$ within the set of designated switchable lines ${E}_{\text{sw}}\subseteq{E}$. For each switchable line, the state variable is defined as
$s_{ij}\in\{0,1\},\quad\forall(i,j)\in{E}_{\text{sw}}.$ (5)
Load curtailment is modeled as a continuous reduction in the active power demand $P_{i}^{\text{load}}$ at buses belonging to the set of curtailable loads ${N}_{\text{L}}\subseteq{N}$. The reduction is parameterized by a curtailment fraction $\gamma_{i}$, resulting in a new load value given by
$P_{i}^{\text{new\_load}}=P_{i}^{\text{orig\_load}}\cdot(1-\gamma_{i}),$ (6)
$0\leq\gamma_{i}\leq\gamma_{\max},\quad\forall i\in{N}_{\text{L}}.$ (7)
Battery placement and dispatch involve two coupled decisions: a discrete choice of where to place batteries, and a continuous choice of how much active and reactive power to inject or absorb. Placement is modeled as a binary decision variable $b_{i}$ for each bus, constrained by the total number of available batteries $B_{\text{total}}$:
$\sum_{i\in{N}}b_{i}\leq B_{\text{total}},\quad\text{with }b_{i}\in\{0,1\}.$ (8)
Once placed, each battery can dispatch active power $P_{B,i}$ and reactive power $Q_{B,i}$, subject to its maximum power rating $S_{B}^{\max}$. Importantly, dispatch is only allowed if a battery is installed at the corresponding bus ($b_{i}=1$). The coupling between placement and operation is enforced as
$P_{B,i}^{2}+Q_{B,i}^{2}\leq(b_{i}\cdot S_{B}^{\max})^{2},$ (9)
$-P_{B}^{\max}\leq P_{B,i}\leq P_{B}^{\max},$ (10)
$-Q_{B}^{\max}\leq Q_{B,i}\leq Q_{B}^{\max}.$ (11)
This formulation guarantees that batteries are both allocated and dispatched within their physical limits, ensuring they can effectively support grid operations.
@@ III-D1 LLMs as Reasoning Engines for Grid Management
Modern LLMs function as powerful reasoning engines. In the context of power systems, they offer several key advantages over traditional algorithms. First, LLMs bring semantic understanding, as they can process and interpret not just raw numerical data, but also heterogeneous sources of information such as network topology descriptions, operational logs, and maintenance reports. This capability enables a more holistic understanding of grid states. Second, they enable contextual reasoning by incorporating high-level domain knowledge and operational constraints, such as prioritizing critical loads or avoiding switching near vulnerable points, which are difficult to encode in conventional optimization models. Finally, recent advancements allow LLMs to operate in a tool-augmented fashion, delegating complex numerical calculations to validated power flow solvers. This integration grounds their high-level reasoning in the physical realities and precise operational constraints of the grid.
@@ III-D2 Multi-Agent Systems for Decomposed Problem Solving
A single monolithic system struggles to manage the entire scope of violation resolution. Agentic AI addresses this limitation by decomposing the problem into specialized sub-tasks, each handled by a dedicated agent. This approach provides several advantages. Different agents can be designed to develop specialized expertise, excelling at specific roles such as topology analysis, action planning, solution validation, or results summarization. In addition, agents can collaborate iteratively, with a Planner agent proposing an action and a Validator agent simulating its impact to provide feedback. This iterative refinement process allows strategies to improve progressively. Finally, the modular structure offers scalability and interpretability: it adapts to grids of varying complexity while providing transparent and auditable reasoning chains through explicit agent interactions.
The multi-agent system can be formulated as
$W=(A,F,C),$ (12)
where $W$ represents the workflow, $A=\{a_{1},a_{2},\ldots,a_{m}\}$ is the set of agents, $F$ defines the flow relationships between agents, and $C$ specifies conditional branching logic. The workflow state $S_{t}$ at iteration $t$ captures the complete system state, including network configuration, detected violations, executed actions, and intermediate results. State transitions are governed by
$S_{t+1}=\mathcal{T}(S_{t},a_{i},\theta),$ (13)
where $\mathcal{T}$ is the transition function, $a_{i}$ is the executing agent, and $\theta$ represents the agent’s parameters and decision logic. This formulation enables systematic exploration of the action space while maintaining solution quality and computational efficiency through intelligent agent coordination.
@@ IV System Architecture and Methodology
The Grid-Agent framework is an autonomous multi-agent system designed to resolve power grid violations by integrating the semantic reasoning of Large Language Models (LLMs) with the numerical precision of validated power system simulations. The methodology is founded on a state-driven workflow where specialized agents collaborate to plan, execute, and validate control actions in a secure, sandboxed environment. This section details the four core pillars of this methodology: the multi-agent workflow, the LLM-guided action planning process, the adaptive network representation scheme, and the integrated safety mechanisms.
@@ IV-A State-Driven Multi-Agent Workflow
The framework’s operation is governed by a cyclic, state-driven workflow, depicted in Fig. 1, which is orchestrated by five specialized agents. The process begins with the Topology Agent, which initializes the resolution process by parsing the grid’s topology and operational data to establish the initial network state and identify existing violations. This foundational context is then passed to the Planner Agent, the core reasoning engine, which leverages its LLM-based intelligence to formulate a strategic, multi-step action plan designed to cohesively resolve the identified issues. Once a plan is formulated, the Executor Agent acts as the interface to the numerical simulation, translating the Planner’s abstract directives into concrete API calls for a power flow solver on a sandboxed copy of the network. Following execution, the Validator Agent ensures the efficacy and safety of the plan by running a new analysis to assess whether violations were resolved without introducing new instabilities. Based on this outcome, it either approves the changes and terminates the workflow or reverts the changes and signals the need for a new planning cycle, ensuring monotonic progress. Finally, upon successful resolution, the Summarizer Agent facilitates explainability by generating a human-readable summary of the solution and structures the entire process into a data entry for continuous learning datasets.
@@ IV-B LLM-Guided Coordinated Action Planning
The central task of finding an optimal action sequence is navigated by the Planner Agent. Unlike traditional optimizers, Grid-Agent uses the LLM as a tool-augmented reasoning engine. This is achieved through a comprehensive system prompt that structures the LLM’s task, defining five key elements:
@@
1) Role Definition: The LLM is instructed to act as an expert power system operator, grounding its responses in domain-specific best practices.
2) State Context: It receives a structured representation of the current network state, including topology, component statuses, and a detailed list of all active violations.
3) Action Space and Constraints: The prompt explicitly defines the set of available actions (e.g., update_switch_status, add_battery, curtail_load) and their associated operational constraints (e.g., maximum number of deployable batteries, available curtailable loads).
4) Strategic Guidance: The LLM is guided by a prioritized action policy that mirrors industry best practices: (1) Topology Reconfiguration, (2) Battery Deployment/Dispatch, and (3) Load Curtailment. This constrains the search space and aligns the LLM’s reasoning with established operational logic.
5) Output Schema: To ensure reliable machine integration, the LLM is required to respond exclusively with a machine-parsable list of tool calls in a structured format (e.g., JSON). This eliminates ambiguous natural language and enables direct interfacing with the Executor Agent.
@@
IV-C Adaptive Multi-Scale Network Representation
IV-D Safety, Validation, and Rollback Mechanisms
Operational safety in a critical infrastructure context is paramount, and Grid-Agent incorporates a multi-layered validation strategy to ensure system stability and reliability. The first layer is a Proactive Execution Check, where the Executor Agent performs a preliminary validation before applying any action in the sandbox. This check uses the power flow solver to ensure the proposed API call is syntactically correct and will not cause an immediate simulation convergence failure, acting as a first line of defense. Following execution, the Validator Agent conducts a Post-Hoc State Assessment, a comprehensive evaluation to verify not only that target violations are resolved but also that no new violations have been introduced elsewhere in the network. Finally, to provide a Monotonic Progress Guarantee, an automated rollback mechanism is triggered if the Validator determines that a plan has failed to improve or has worsened the grid state. This mechanism reverts the sandboxed network to its state before the failed attempt, ensuring the system only progresses towards a verifiably better solution. This defense-in-depth approach, combining proactive checks with comprehensive post-hoc validation and automated rollbacks, provides the essential safety guarantees for deploying an AI-driven system in a mission-critical environment.
IV-E Continuous Learning Framework